# CBF-RL: Safety Filtering Reinforcement Learning in Training with Control Barrier Functions

### v2

@@ III Dual Approach to CBF-RL
Suppose $h:\mathbb{R}^{n}\to\mathbb{R}$ is a $C^{1}$ CBF for the continuous-time single integrator $\dot{q}=v$, where $q,v\in\mathbb{R}^{n}$. Let $k_{s}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be any safe, locally Lipschitz controller for the continuous-time integrator, satisfying
$\nabla h(\mathbf{q})^{\top}k_{s}(\mathbf{q})\geq-\alpha h(\mathbf{q}),\;\forall\mathbf{q}\in\mathbb{R}^{n},$ (11)
for some $\alpha>0$. Let $\mathbf{f}_{\mathrm{\Delta t}}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{n},(\mathbf{q},\mathbf{v})\mapsto\mathbf{q}+\mathrm{\Delta t}\,\mathbf{v}$ be the Euler discretization of the single integrator with time step $\mathrm{\Delta t}>0$. There exists a continuous function $R:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ such that $\forall\mathbf{q}\in\mathbb{R}^{n}$, $\lim_{\left\lVert\mathbf{w}\right\rVert\to 0}\frac{R(\mathbf{q},\mathbf{w})}{\left\lVert\mathbf{w}\right\rVert}=0$ and
@@
By [36, Chapter 5.2 (2)], there exists a continuous function $R:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfying $\lim_{\left\lVert\mathbf{w}\right\rVert\to 0}\frac{R(\mathbf{q},\mathbf{w})}{\left\lVert\mathbf{w}\right\rVert}=0$ and $h(\mathbf{q}+\mathbf{w})=h(\mathbf{q})+\nabla h(\mathbf{q})^{\top}\mathbf{w}+R(\mathbf{q},\mathbf{w})\;\forall\mathbf{q},\mathbf{w}\in\mathbb{R}^{n}$. Taking $\mathbf{w}=\mathrm{\Delta t}\,k_{s}(\mathbf{q})$, we calculate $h(\mathbf{q}+\mathrm{\Delta t}\,k_{s}(\mathbf{q}))$ as
@@
Consider the setting of Lemma 1. Suppose there exists a compact, forward-invariant set $K\subseteq\mathbb{R}^{n}$ for the discrete-time integrator dynamics $\mathbf{q}_{k+1}=\mathbf{f}_{\mathrm{\Delta t}}(\mathbf{q}_{k},k_{s}(\mathbf{q}_{k}))$. Then, $\forall\mathrm{\Delta t}>0$, $\mu(\mathrm{\Delta t})=\sup_{\mathbf{q}\in K}|R(\mathbf{q},\mathrm{\Delta t}\,k_{s}(\mathbf{q}))|<+\infty$ and $\lim_{\mathrm{\Delta t}\to 0}\mu(\mathrm{\Delta t})/\mathrm{\Delta t}=0$. Further, provided $(1-\mathrm{\Delta t}\,\alpha)\in[0,1)$, for $\mathbf{q}_{0}\in K$,
@@
The established relationship means that with small enough $\Delta t$, as is the norm with physical simulators oriented towards RL, we can apply continuous-time CBF tools directly to discrete-time RL environments; thus we provide the two core parts of CBF-RL shown in Fig. 2:
Safety-filtering during training At training time, the policy would generate desired actions that are not necessarily safe $\mathbf{v}_{k}^{\mathrm{policy}}$ at step $k$. A safety filter is then applied to enforce safe behaviors and guide the system to ’learn’ the safety filter as part of the natural closed-loop dynamics. As we have proven the relationship between the continuous-time and discrete-time CBF conditions, typically, we can achieve safety filtering through a CBF-QP as follows [9]:
@@
However, solving QPs for every step of RL training is not preferable in the massively-parallel environments such as IsaacLab [22], instead, we employ the explicit solution:
@@
$r_{{\mathrm{cbf}}}(\mathbf{q}_{k},\mathbf{v}_{k})=$ $\max\!\left(\left(\mathbf{a}_{k}^{\top}\mathbf{v}^{\mathrm{policy}}_{k}-b_{k}\right),\,0\right)$ (22)
@@
5:   $\mathbf{q}^{\mathrm{policy}}_{k+1}\leftarrow\pi_{\theta}(\mathbf{o}_{k})$
6:   $\mathbf{v}^{\mathrm{policy}}_{k}\leftarrow\frac{(\mathbf{q}^{\mathrm{policy}}_{k+1}-\mathbf{q}_{k})}{\Delta t}$
@@
9:   if $c\geq 0$ then
10:     $\mathbf{v}^{\mathrm{safe}}_{k}\leftarrow\mathbf{v}^{\mathrm{policy}}_{k}$,
11:   else
12:     $\mathbf{v}^{\mathrm{safe}}_{k}\leftarrow\mathbf{v}^{\mathrm{policy}}_{k}+\dfrac{(b_{k}-\mathbf{a}_{k}^{\top}v^{\mathrm{policy}}_{k})\,\mathbf{a}_{k}^{\!\top}}{\|\mathbf{a}_{k}\|^{2}}$
13:   end if
14:   $\mathbf{q}^{\mathrm{safe}}_{k+1}\leftarrow\mathbf{q}_{k}+\Delta t\,\mathbf{v}^{\mathrm{safe}}_{k}$
15:   $\mathbf{q}^{\mathrm{env}}_{k+1},\mathbf{o}_{k+1}\leftarrow\textsc{environment\_update}(\mathbf{q}^{\mathrm{safe}}_{k+1})$
16:   $r_{\mathrm{cbf}}\leftarrow w\cdot[\max\!\left(0,\,\left(\mathbf{a}_{k}^{\top}\mathbf{v}^{\mathrm{policy}}_{k}-b_{k}\right)\right)+\exp\!\left(-\tfrac{\|\mathbf{v}^{\mathrm{policy}}-\mathbf{v}^{\mathrm{safe}}\|^{2}}{\sigma^{2}}\right)-1]$
17:   $r\leftarrow R(\mathbf{q}^{\mathrm{env}}_{k+1})+r_{\mathrm{cbf}}$
18:   Store transition: $(\mathbf{q}_{k},\mathbf{o}_{k},\mathbf{q}^{\mathrm{policy}}_{k+1},\mathbf{q}^{\mathrm{safe}}_{k+1},\mathbf{q}^{\mathrm{env}}_{k+1},r)$
19:  end for
20:  Update policy parameters $\theta\leftarrow\eta\nabla_{\theta}\mathcal{L}(\theta,r)$
21:end for
Intuitively, this reward penalizes actions whenever the safety filter is activated, and also incentivizes the model to take actions as close to the safe actions as possible to reduce the intervention of the filter. The whole algorithm is as shown in Alg. 1.
Single Integrator Example To demonstrate our proposed approach and analyze the effect of each component of CBF-RL, we perform extensive ablation studies on a single integrator navigation task. The agent is placed into a 2D world with obstacles and tasked with going to a specified goal. The starting, obstacle, and goal locations are all randomized, and extra care is taken such that the agent always initializes in a safe state following [13]. The single integrator has dynamics $\mathbf{q}_{k+1}=\mathbf{q}_{k}+\mathbf{v}_{k}\Delta t$ where $\mathbf{q}=(x,y)$ is the robot position, $\mathbf{v}_{k}$ is the velocity of the agent, and $\Delta t$ is the timestep size. The safety barrier function $h$ is defined as
@@
Ablation To validate our method, we train 4 variants (Dual, Reward-only, Filter-only, Nominal) and test 12 variants following Table III for 1500 steps with 4096 parallel environments. We also evaluate policies trained with filtered action then deployed without a runtime filter (rt. filt.). We observe that the Dual approach achieves higher rewards while remaining safe throughout training. Over 1000 random test environments, the Dual policy performs well with or without a runtime filter, whereas Filter Only performs well only with an active safety filter and degrades markedly without it.
@@
$r_{\mathrm{cbf}}$ $\begin{aligned} &100\cdot\Big(\max(\nabla h(\mathbf{q})^{\top}\mathbf{v}+\alpha h(\mathbf{q}),\,0)\\ &+\exp\!\big(-\tfrac{\|\mathbf{v}_{\mathrm{policy}}-\mathbf{v}_{\mathrm{safe}}\|^{2}}{0.5^{2}}\big)-1\Big)\cdot\mathbf{1}(\text{active})\end{aligned}$
@@ IV CBF-RL for Safe Humanoid Locomotion
In the following section, we present two different use cases of CBF-RL in humanoid locomotion to show the generality and performance of our method. Each uses a different set of user-specified reduced-order coordinates and a valid CBF for the resulting reduced-order dynamics. The nominal rewards and observations (history length 5) follow [37]. Note that for blind stair climbing, we employ the asymmetric actor-critic method [38] and provide a height scan of size 1m $\times$ 1.5m with a resolution of 0.1m to the critic observation with a history length of 1. We use a MLP policy with 3 hidden layers of $[512,256,128]$ neurons running at 5050Hz, outputting the joint position setpoints of the 12-DoF lower body for the Unitree G1. We train in IsaacLab with 4096 environments of $\Delta t=0.005s$ up to 20,000 steps and perform zero-shot hardware transfer experiments to verify our approach.
@@ IV-A Task Definition
$r_{\mathrm{next\,stair\,cbf}}=$ $\max\!\left(\;\big(-\mathbf{J}^{\mathrm{sw}}_{x}(\mathbf{q})\big)\mathbf{v}+\alpha\,h(\mathbf{q}),0\right)$
@@ IV-B Hardware Experiments
Obstacle Course The obstacle course comprises 2 parts: the first part is the obstacle avoidance task where the robot has to prevent itself from colliding into the obstacles even if the velocity command intends so, and the second part comprises stairs constructed out of wooden pallets of riser height 0.14m and tread depth 0.3m. During execution, the robot locates the obstacles approximated as cylinders using the ZED 2 RGB-D camera through point cloud clustering. As seen from the $h$ values as in Fig. 5, the robot modulates its own velocity to avoid the obstacle even when the velocity command pushes it to do so. However, it has no terrain perception and only uses proprioception for stair climbing. Despite this, we observe that the robot uses proprioception to determine when to climb and how high to lift its feet, successfully climbing the wooden stairs.
Stair Climbing Robustness In order to further test the robustness of our method, we experiment both indoors and outdoors. During indoor experiments, we perform continuous runs up and down stairs, and also test the performance of the policy on high stairs, with the dual-trained policy able to climb up stairs 0.3m high and the policy trained without the CBF modifications unable to do so, as shown in Fig. 6. Here we note that the robot is able to gauge the depth and height of the stairs through proprioception and adjust its footsteps accordingly. For outdoor experiments, we test the dual-trained policy on concrete-poured stairs of different roughness and sizes, with rougher stairs of riser height 0.14m / tread depth 0.33m and smoother stairs of riser height 0.15m / tread depth 0.4m respectively, as shown in Fig. 7, where the robot could adjust its center of mass by modulating the torso pitch angle to account for deeper and higher stairs.
@@ V Conclusion
This paper introduced CBF-RL, a lightweight dual approach to inject safety into learning by combining training-time CBF safety filtering with reward design, leading to policies that internalize safety and operate without a runtime filter, demonstrating its effectiveness through simulated and real-world experiments. We also provided a detailed theoretical analysis rationalizing the use of continuous-time CBF conditions in discrete-time RL simulation training. Looking forward, we plan to incorporate automated barrier discovery, perception-based barriers, and extend the application of CBF-RL beyond locomotion to whole-body loco-manipulation that would address broader humanoid capabilities.

### v4

@@ III Dual Approach to CBF-RL
Suppose $h:\mathbb{R}^{n_{q}}\to\mathbb{R}$ is a $C^{1}$ CBF for the continuous-time single integrator $\dot{q}=v$, where $q,v\in\mathbb{R}^{n_{q}}$. Let $k_{s}:\mathbb{R}^{n_{q}}\to\mathbb{R}^{n_{q}}$ be any safe, locally Lipschitz controller for the continuous-time integrator, satisfying
$\nabla h(\mathbf{q})^{\top}k_{s}(\mathbf{q})\geq-\alpha h(\mathbf{q}),\;\forall\mathbf{q}\in\mathbb{R}^{n_{q}},$ (11)
for some $\alpha>0$. Let $\mathbf{f}_{\mathrm{\Delta t}}:\mathbb{R}^{n_{q}}\times\mathbb{R}^{n_{q}}\to\mathbb{R}^{n_{q}},(\mathbf{q},\mathbf{v})\mapsto\mathbf{q}+\mathrm{\Delta t}\,\mathbf{v}$ be the Euler discretization of the single integrator with time step $\mathrm{\Delta t}>0$. There exists a continuous function $R:\mathbb{R}^{n_{q}}\times\mathbb{R}^{n_{q}}\to\mathbb{R}$ such that $\forall\mathbf{q}\in\mathbb{R}^{n_{q}}$, $\lim_{\left\lVert\mathbf{w}\right\rVert\to 0}\frac{R(\mathbf{q},\mathbf{w})}{\left\lVert\mathbf{w}\right\rVert}=0$ and
@@
By [pugh2002real, Chapter 5.2 (2)], there exists a continuous function $R:\mathbb{R}^{n_{q}}\times\mathbb{R}^{n_{q}}\to\mathbb{R}$ satisfying $\lim_{\left\lVert\mathbf{w}\right\rVert\to 0}\frac{R(\mathbf{q},\mathbf{w})}{\left\lVert\mathbf{w}\right\rVert}=0$ and $h(\mathbf{q}+\mathbf{w})=h(\mathbf{q})+\nabla h(\mathbf{q})^{\top}\mathbf{w}+R(\mathbf{q},\mathbf{w})\;\forall\mathbf{q},\mathbf{w}\in\mathbb{R}^{n_{q}}$. Taking $\mathbf{w}=\mathrm{\Delta t}\,k_{s}(\mathbf{q})$, we calculate $h(\mathbf{q}+\mathrm{\Delta t}\,k_{s}(\mathbf{q}))$ as
@@
Consider the setting of Lemma 1. Suppose there exists a compact, forward-invariant set $K\subseteq\mathbb{R}^{n_{q}}$ for the discrete-time integrator dynamics $\mathbf{q}_{k+1}=\mathbf{f}_{\mathrm{\Delta t}}(\mathbf{q}_{k},k_{s}(\mathbf{q}_{k}))$. Then, $\forall\mathrm{\Delta t}>0$, $\mu(\mathrm{\Delta t})=\sup_{\mathbf{q}\in K}|R(\mathbf{q},\mathrm{\Delta t}\,k_{s}(\mathbf{q}))|<+\infty$ and $\lim_{\mathrm{\Delta t}\to 0}\mu(\mathrm{\Delta t})/\mathrm{\Delta t}=0$. Further, provided $(1-\mathrm{\Delta t}\,\alpha)\in[0,1)$, for $\mathbf{q}_{0}\in K$,
@@
The above means that with small enough $\Delta t$ (typically $\Delta t\leq 0.01$s, governed by the stability limits and solve time of the physics engine), continuous-time CBF tools can be directly applied to discrete-time RL environments. We provide the two core parts of CBF-RL shown in Fig. 2:
Safety-filtering during training At training time, the policy would generate desired actions that are not necessarily safe $\mathbf{v}_{k}^{\mathrm{policy}}$ at step $k$. A safety filter is then applied to enforce safe behaviors and guide the system to ’learn’ the safety filter as part of the natural closed-loop dynamics. As we have proven the relationship between the continuous-time and discrete-time CBF conditions, we replace the typically nonlinear DTCBF constraint with the continuous-time CBF first-order inequality. This reduces the safety filtering problem to a single linear-constraint quadratic program (QP) [ames2019control]:
@@
While this QP provides a safe control input, solving an optimization program numerically at every step of RL training is computationally undesirable, especially in massively parallel environments such as IsaacLab [mittal2023orbit]. Fortunately, since the QP has a single linear constraint, it can be solved analytically in closed-form:
@@
$r_{{\mathrm{cbf}}}(\mathbf{q}_{k},\mathbf{v}_{k})=$ $\min\!\left(\mathbf{a}_{k}^{\top}\mathbf{v}^{\mathrm{policy}}_{k}-b_{k},\,0\right)$ (22)
@@
Intuitively, this reward penalizes actions whenever the safety filter is activated, and also incentivizes the model to take actions as close to the safe actions as possible to reduce the intervention of the filter. Because the penalty term $\exp(\dots)-1$ is strictly lower-bounded by 1-1, it provides a smooth learning signal without causing unbounded instability with respect to the primary task rewards. The whole algorithm is as shown in Alg. 1.
@@
5:   $\mathbf{v}^{\mathrm{policy}}_{k}\leftarrow\pi_{\theta}(\mathbf{o}_{k})$
6:   $\mathbf{q}^{\mathrm{policy}}_{k+1}\leftarrow\mathbf{q}_{k}+\Delta t\,\mathbf{v}^{\mathrm{policy}}_{k}$
@@
9:   if $c\geq 0$ then $\mathbf{v}^{\mathrm{safe}}_{k}\leftarrow\mathbf{v}^{\mathrm{policy}}_{k}$,
10:   else $\mathbf{v}^{\mathrm{safe}}_{k}\leftarrow\mathbf{v}^{\mathrm{policy}}_{k}+\dfrac{(b_{k}-\mathbf{a}_{k}^{\top}v^{\mathrm{policy}}_{k})\,\mathbf{a}_{k}^{\!\top}}{\|\mathbf{a}_{k}\|^{2}}$
11:   $\mathbf{q}^{\mathrm{safe}}_{k+1}\leftarrow\mathbf{q}_{k}+\Delta t\,\mathbf{v}^{\mathrm{safe}}_{k}$
12:   $\mathbf{q}^{\mathrm{env}}_{k+1},\mathbf{o}_{k+1}\leftarrow\textsc{environment\_update}(\mathbf{q}^{\mathrm{safe}}_{k+1})$
13:   $r\leftarrow w\cdot[\min\!\left(\mathbf{a}_{k}^{\top}\mathbf{v}^{\mathrm{policy}}_{k}-b_{k},0\right)+\exp\!\left(-\tfrac{\|\mathbf{v}^{\mathrm{policy}}-\mathbf{v}^{\mathrm{safe}}\|^{2}}{\sigma^{2}}\right)-1]+R(\mathbf{q}^{\mathrm{env}}_{k+1})$
14:   Store transition: $(\mathbf{q}_{k},\mathbf{o}_{k},\mathbf{q}^{\mathrm{policy}}_{k+1},\mathbf{q}^{\mathrm{safe}}_{k+1},\mathbf{q}^{\mathrm{env}}_{k+1},r)$
15:  end for
16:  Update policy parameters $\theta\leftarrow\eta\nabla_{\theta}\mathcal{L}(\theta,r)$
17:end for
Single Integrator Example To demonstrate our proposed approach and analyze the effect of each component of CBF-RL, we perform extensive ablation studies on a single integrator navigation task. The agent is placed into a 2D world with obstacles and tasked with going to a specified goal. The starting, obstacle, and goal locations are all randomized, and extra care is taken such that the agent always initializes in a safe state following [bejarano_safety_2025]. The single integrator has dynamics $\mathbf{q}_{k+1}=\mathbf{q}_{k}+\mathbf{v}_{k}\Delta t$ where $\mathbf{q}=(x,y)$ is the robot position, $\mathbf{v}_{k}$ is the velocity of the agent, and $\Delta t$ is the timestep size. The safety barrier function $h$ is defined as
@@
Ablation To validate our method, we train 4 variants (Dual, Reward-only, Filter-only, Nominal) and test 12 variants following Table III for 1500 steps with 4096 parallel environments. We also evaluate policies trained with filtered action then deployed without a runtime filter (rt. filt.). The training progress can be seen in Fig. 4, where we observe that the Dual and Filter-only approaches achieve rapid convergence while remaining safe throughout training. Furthermore, as illustrated by the trajectory comparisons in Fig. 4 as an example from over 1000 random test environments, the Dual approach is able to reach the goal in both domain randomized and non-domain randomized settings, even without a runtime filter, while the other methods fail to do so. Notably, the Filter Only approach performs well only with an active safety filter and degrades markedly without it.
@@
$r_{\mathrm{cbf}}$ $\begin{aligned} &100\cdot\Big(\min(\nabla h(\mathbf{q})^{\top}\mathbf{v}+\alpha h(\mathbf{q}),\,0)\\ &+\exp\!\big(-\tfrac{\|\mathbf{v}_{\mathrm{policy}}-\mathbf{v}_{\mathrm{safe}}\|^{2}}{0.5^{2}}\big)-1\Big)\cdot\mathbf{1}(\text{active})\end{aligned}$
@@ IV CBF-RL for Safe Humanoid Locomotion
In the following section, we present two different use cases of CBF-RL in humanoid locomotion to show the generality and performance of our method. Each uses a different set of user-specified reduced-order coordinates and a valid CBF for the resulting reduced-order dynamics. The nominal rewards and observations (history length 5) follow [zakka2025mujoco]. Note that for blind stair climbing, we employ the asymmetric actor-critic method [pinto2017asymmetric] and provide a height scan of size 1m $\times$ 1.5m with a resolution of 0.1m to the critic observation with a history length of 1. We use an MLP policy with 3 hidden layers of $[512,256,128]$ neurons running at 5050Hz, outputting the joint position setpoints of the 12-DoF lower body for the Unitree G1. We train in IsaacLab with 4096 environments with $\Delta t=0.005$s. Each episode lasts for a maximum of 20,000 steps on the NVIDIA RTX 4090 GPU and perform zero-shot hardware transfer experiments to verify our approach.
@@ IV-A Task Definition
$r_{\mathrm{next\,stair\,cbf}}=$ $\min\!\left(\;\big(-\mathbf{J}^{\mathrm{sw}}_{x}(\mathbf{q})\big)\mathbf{v}+\alpha\,h(\mathbf{q}),0\right)$
@@
Hardware Deployment Considerations The stochasticity of real-world state estimation hinders the deployment of runtime explicit filters. CBF-RL is able to internalize safety by learning to map noisy observations to inherently safe actions, thus reducing the reliance of such a filter.
@@ IV-B Hardware Experiments
Obstacle Course The obstacle course comprises 2 parts: the first part is the obstacle avoidance task, where the robot has to prevent itself from colliding with the obstacles, even if the velocity command intends otherwise; the second part comprises stairs constructed out of wooden pallets with a riser height of 0.14m and tread depth of 0.3m. During execution, the robot locates the obstacles approximated as cylinders using the ZED 2 RGB-D camera through point cloud clustering. As seen from the $h$ values in Fig. 5, the robot modulates its own velocity to avoid the obstacle, even when the velocity command prompts it to do so. However, it has no terrain perception and only uses proprioception for stair climbing. Despite this, we observe that the robot uses proprioception to determine when to climb and how high to lift its feet, successfully climbing the wooden stairs.
Stair Climbing Robustness In order to further test the robustness of our method, we experiment both indoors and outdoors. During indoor experiments, we perform continuous runs up and down stairs and also test the performance of the policy on high stairs, with the dual-trained policy able to climb up stairs 0.3m high and the policy trained without the CBF modifications unable to do so, as shown in Fig. 6. Here we note that the robot is able to gage the depth and height of the stairs through proprioception and adjust its footsteps accordingly. For outdoor experiments, we test the dual-trained policy on concrete-poured stairs of different roughness and sizes, with rougher stairs of riser height 0.14m / tread depth 0.33m and smoother stairs of riser height 0.15m / tread depth 0.4m respectively, as shown in Fig. 7, where the robot could adjust its center of mass by modulating the torso pitch angle to account for deeper and higher stairs.
@@ V Conclusion
This paper introduces CBF-RL, a lightweight dual approach to inject safety into learning by combining training-time CBF safety filtering with reward design, leading to policies that internalize safety and operate without a runtime filter, demonstrating its effectiveness through simulated and real-world experiments. Looking forward, we plan to incorporate automated barrier discovery, perception-based barriers, and extend the application of CBF-RL beyond locomotion to whole-body loco-manipulation, which will address broader humanoid capabilities.
Acknowledgment During the preparation of this work, the authors used AI-assisted capabilities (Claude, ChatGPT) for grammar and editing enhancements of the text, as well as for debugging software code related to the experiments. The authors have responsibly reviewed and verified all AI-generated content to ensure correctness.